OpenStudy (anonymous):

find the derivative of the following function y=x^(cos4x)

sam (.sam.):

\[\large y=x^{\cos(4x)} \] Natural log both sides, \[ \large \ln(y)=\cos(4x)\ln(x)\] Implicit differentiation it

sam (.sam.):

Do you understand?

OpenStudy (anonymous):